# The significance of measuring cosmological time dilation in the Dark Energy Survey Supernova Program

### v2

@@ Abstract
In the context of the dispersion relation $c=\lambda\nu$ and considering an expanding universe where the observed wavelength today is redshifted from the emitted wavelength by $\lambda_{0}=\lambda_{\text{emit}}(1+z)$, to keep $c$ constant, it must be that $\nu_{0}=\nu_{\text{emit}}/(1+z)$. However, although the theory for wavelength in the RW metric includes the cosmological redshift, the same is not simply deduced for frequency (the inverse of time). Instead, cosmological time dilation $T_{0}=T_{\text{emit}}(1+z)$ is an additional assumption made to uphold the hypothesis of constant speed of light rather than a relation directly derived from the RW metric. Therefore, verifying cosmological time dilation observationally is crucial. The most recent data employing supernovae for this purpose was released recently by the Dark Energy Survey. Results from the i-band specifically support variations in the speed of light within 1-$\sigma$. We used these observations to investigate variations in various physical quantities, including $c$ and $G$, using the minimally extended varying speed of light model. The speed of light was $0.4$% to $2.2$% slower, and Newton’s constant may have decreased by $1.7$% to $8.4$% compared to their current values at redshift $2$. These findings, consistent with previous studies, hint at resolving tensions between different $\Lambda$CDM cosmological backgrounds but are not yet conclusive evidence of a varying speed of light, as the full-band data aligns with standard model cosmology.
@@ II Methods
II.1 Cosmological time dialtion from standard clocks
@@
A standard clock with a measured timescale is necessary to look for cosmological time dilation. However, in the case of other cosmic objects like quasars and GRBs, the complexity of the physical processes causing their variability makes it difficult to use them as standard clocks. Nevertheless, the standard clock method using QSos and GRBs has been consistently observed.
@@
In analyzing $247$ GRBs, intrinsic scatter greater than $100$% complicates establishing a clear relationship between $T_{50}/T_{90}$ and redshift Singh:2021jgr . The power-law index $B$ differs from the expected value of one within $1$-$\sigma$. By focusing on long GRBs and using geometric means to reduce outlier effects, a tighter scatter is observed. For $T_{50}$, the intrinsic scatter is less than $60$%, and for $T_{90}$, it is consistent with $27$% within a $68$% confidence level, indicating the best-fit scatter value exceeds $100$%. The power-law exponent $B$ aligns with a cosmological signature within 1-$\sigma$ for both $T_{50}$ and $T_{90}$. This analysis supports the presence of CTD, with the geometric mean providing the most consistent results ($n\equiv 1+\beta=1.18^{+0.26}_{-0.36}$).
@@ III Results
The DES evaluated SNeIa LCs detected in the $g$, $r$, $i$, and $z$ bands using the reference-scaling method with precision to obtain CTD. By analyzing data from all the bands combined, 1504 distinct SNeIa were included in the analysis. The analysis yielded a reduced chi-square value of $1.441$ and a time dilation relation of $(1+z)^{1.003\pm 0.005}$, which aligns with the SMC within the 1-sigma confidence level.
In contrast, 1465 SNe satisfied the quality criterion when the analysis was limited to the $i$ band. The decreased chi-square value for these SNeIa was $0.788$. $(1+z)^{n}$, with $n=0.988\pm 0.008$, provided the best match for the time dilation. This data suggests that the CTD may be affected by the VSL, as it deviates from the SMC within the $1$-$\sigma$ confidence interval.
@@ IV Conclusion and Summary
Using the recently released dark energy survey supernova program’s $i$ band results, we applied the minimally extended varying speed of light model to investigate the cosmological evolution of the speed of light and Newton’s constant within the 1-$\sigma$ confidence levels. Our analysis indicates that the speed of light was between $0.4$% and $2.2$% slower at redshift $2$ compared to its current value. Furthermore, the gravitational constant may have decreased by $1.7$% to $8.4$% at $z=2$ compared to its present value.
These findings align with previous studies suggesting that Newton’s constant $G(z)$ decreasing with redshift could resolve the observed tension between the Planck15151515 best-fit $\Lambda$CDM cosmological background and the $\Lambda$CDM background favored by growth $f\sigma_{8}$ and weak lensing data. However, the full-band cosmological time dilation data from the dark energy survey supernova program matches the results from the standard model cosmology, indicating that these findings alone are not yet conclusive evidence of a varying speed of light.
It is important to note that the redshift $z$ values used in this data do not account for the additional effects predicted by the minimally extended varying speed of light model. For a consistent interpretation, the redshift values predicted by the minimally extended varying speed of light model should be used when analyzing the data. It necessitates recalculating the redshift for each dataset in the dark energy survey supernova program to compute the cosmological time dilation more accurately. While this task is beyond the scope of this paper, we plan to undertake it in future research.

### v3

@@ Abstract
In the context of the dispersion relation $c=\lambda\nu$ and considering an expanding universe where the observed wavelength today is redshifted from the emitted wavelength by $\lambda_{0}=\lambda_{\text{emit}}(1+z)$, to keep $c$ constant, it must be that $\nu_{0}=\nu_{\text{emit}}/(1+z)$. However, although the theory for wavelength in the RW metric includes the cosmological redshift, the same is not simply deduced for frequency (the inverse of time). Instead, cosmological time dilation $T_{0}=T_{\text{emit}}(1+z)$ is an additional assumption made to uphold the hypothesis of constant speed of light rather than a relation directly derived from the RW metric. Therefore, verifying cosmological time dilation observationally is crucial. The most recent data employing supernovae for this purpose was released recently by the Dark Energy Survey. Results from the i-band specifically support variations in the speed of light within 1-$\sigma$. We used these observations to investigate variations in various physical quantities, including $c$ and $G$, using the minimally extended varying speed of light model. The speed of light was $0.4$% to $2.2$% slower, and Newton’s constant may have decreased by $1.7$% to $8.4$% compared to their current values at redshift $2$. These findings, consistent with previous studies, hint at resolving tensions between different $\Lambda$CDM cosmological backgrounds but are not yet conclusive evidence of a varying speed of light, as the full-band data aligns with standard model cosmology. However, the data remains valuable for testing variations in fundamental constants over cosmic time. Future analyses, particularly with more refined redshift data, may provide clearer insights into these potential changes.
@@ II Methods
II.1 Cosmological time dilation from standard clocks
@@
A standard clock with a measured timescale is necessary to look for CTD. However, in the case of other cosmic objects like quasars and GRBs, the complexity of the physical processes causing their variability makes it difficult to use them as standard clocks. Nevertheless, the standard clock method using QSos and GRBs has been consistently observed.
@@
In analyzing $247$ GRBs, intrinsic scatter greater than $100$% complicates establishing a clear relationship between $T_{50}/T_{90}$ and redshift Singh:2021jgr . The power-law index $\beta$ differs from the expected value of one within $1$-$\sigma$. By focusing on long GRBs and using geometric means to reduce outlier effects, a tighter scatter is observed. For $T_{50}$, the intrinsic scatter is less than $60$%, and for $T_{90}$, it is consistent with $27$% within a $68$% confidence level, indicating the best-fit scatter value exceeds $100$%. The power-law exponent $\beta$ aligns with a cosmological signature within 1-$\sigma$ for both $T_{50}$ and $T_{90}$. This analysis supports the presence of CTD, with the geometric mean providing the most consistent results ($n\equiv 1+\beta=1.18^{+0.26}_{-0.36}$).
@@ III Results
The DES evaluated SNeIa LCs detected in the $g$, $r$, $i$, and $z$ bands using the reference-scaling method with precision to obtain CTD DES:2024vgg . By analyzing data from all the bands combined, 1504 distinct SNeIa were included in the analysis. The analysis yielded a reduced chi-square value of $1.441$ and a time dilation relation of $(1+z)^{1.003\pm 0.005}$, which aligns with the SMC within the 1-sigma confidence level. However, this result only includes statistical errors. In analyzing supernova light curves, one might need to consider potential systematic effects that could influence the observed data. One such effect is the evolution of the stretch of supernova light curves as a function of redshift. After examining this factor, DES found that the impact is relatively small, with a likely upper limit of around $\sigma^{\textrm{sys}}=0.01$. Therefore, we conduct the following analysis by dividing it into cases with and without this systematic error.
III.1 all-band without systematic error
III.1.1 $i$-band without systematic error
In contrast, 1465 SNe satisfied the quality criterion when the analysis was limited to the $i$ band. The decreased chi-square value for these SNeIa was $0.788$. The DES analyzes $(1+z)^{n}$, with $n=0.988\pm 0.008$, provided the best match for the time dilation. This data suggests that the CTD may be affected by the VSL, as it deviates from the SMC within the $1$-$\sigma$ confidence interval.
@@ III.1.1 $i$-band without systematic error
III.1.2 other bands without systematic error
Table 3 presents the upper bounds for the parameter $b$, which reflects potential variations in the speed of light and Newton’s constant, obtained from different bands of data. The columns list the bands (e..g., $g$, $r$, $i$, $z$, all), the $b$-value ranges, the derived ratio of the speed of light at redshift $z=2$ to its current value $c_{0}$, and the ratio of Newton’s constant at redshift $z=2$ to its current value $G_{0}$. The final column, labeled SMC, indicates whether the results for each band are consistent with the SMC, where the speed of light and Newton’s constant are assumed to be constant. For the $g$, $r$, and $z$ bands, the data supports the SMC, as indicated by an O, while the i band shows inconsistency (marked with an X). The all row aggregates results across all bands, showing consistency with the SMC. The table suggests that the current DES data generally supports the constancy of these fundamental constants, except for the $i$ band.
Bands $b$-value $c(z=2)/c_{0}$ $G(z=2)/G_{0}$ SMC
$g$ $[-0.192\,,0.024]$ $[0.993\,,1.054]$ $[0.974\,,1.235]$ O
$r$ $[-0.048\,,0.016]$ $[0.996\,,1.013]$ $[0.983\,,1.054]$ O
$i$ $[0.016\,,0.080]$ $[0.978\,,0.996]$ $[0.916\,,0.982]$ X
$z$ $[-0.020\,,0.044]$ $[0.988\,,1.006]$ $[0.953\,,1.022]$ O
all $[-0.032\,,0.008]$ $[0.998\,,1.009]$ $[0.991\,,1.036]$ O
Table 3: These are the upper bounds for the range of $b$ obtained from each band and the possible variations in the speed of light and Newton’s constant derived from them. The current DES data shows results consistent with the SMC (i.e., constant speed of light) for all bands except the $i$-band.
  • $g$ band: The $b$-value can vary from $-0.192$ to $0.024$. This indicates possible deviations from the constant speed of light and Newton’s constant. The ratio of the speed of light at $z=2$ to its current value, $c(z=2)/c_{0}$, can vary from $0.993$ to $1.054$, meaning there is a small range for the speed of light to be slightly slower or faster at redshift $z=2$ compared to today. For Newton’s constant, $G(z=2)/G_{0}$, the values range from $0.974$ to $1.235$, suggesting slightly weaker or stronger. The results are consistent with the SMC, as noted by the O.
  • $r$ band: Here, the $b$-value ranges from $-0.048$ to $0.016$. The speed of light ratio can vary from $0.996$ to $1.013$, showing only minimal potential variation from its present value. Similarly, the ratio $G(z=2)/G_{0}$ ranges from $0.983$ to $1.054$, indicating a limited possible deviation in Newton’s constant. Like the $g$ band, the results are consistent with the SMC.
  • $i$ band: The $b$-value for the i band ranges from $0.016$ to $0.080$. For the speed of light, $c(z=2)/c_{0}$ could vary between $0.978$ and $0.996$, implying that the speed of light might have been slower at $z=2$. The ratio $G(z=2)/G_{0}$ ranges from $0.916$ to $0.982$, suggesting that Newton’s constant might also have been weaker. The $i$ band is inconsistent with the SMC, as indicated by the X.
  • $z$ band: The $b$-value in the $z$ band ranges from $-0.020$ to $0.044$. The speed of light ratio can vary between $0.988$ and $1.006$, meaning the speed of light could have been either marginally slower or faster at redshift $2$. The ratio $G(z=2)/G_{0}$ ranges from $0.953$ to $1.022$, suggesting only minor variations in Newton’s constant. The $z$ band is consistent with the SMC.
  • All bands combined: When all bands are combined, the $b$-value ranges from $-0.032$ to $0.008$, providing more constrained upper limits for potential variations. The speed of light ratio lies between $0.998$ and $1.009$, while the ratio of Newton’s constant varies from $0.991$ to $1.036$. This result suggests only slight variations, supporting the idea of constant fundamental constants. It indicates consistency with the SMC for the combined data across all bands.
III.2 all-band with systematic error
III.2.1 $i$-band with systematic error
A possible systematic effect due to the evolution of the stretch of the supernova light curve provides the value $n=0.988\pm 0.008(\textrm{stat})\pm 0.010(\textrm{sys})$ including the maximum error. From this result, we obtain the value of the parameter $b$ of the meVSL model as
$b=0.048\pm 0.051\,.$ (7)
From this value, the values of the speed of light and Newton’s constant at $z=2$ are given as
$0.973\leq\frac{c(z=2)}{c_{0}}\leq 1.001\quad,\quad 0.897\leq\frac{G(z=2)}{G_{0}}\leq 1.003\,,$ (8)
In this case, they are marginally consistent with the values in the standard model.
Refer to caption Refer to caption
Figure 2: The possible evolutions of the speed of light and Newton’s gravitational constant over cosmic time, normalized to their current values when we include the systematic error. Left) The evolution of the speed of light, normalized to its current value, from $z=0$ to $z=2$. Right) The evolution of Newton’s gravitational constant over the same redshift range.
The left panel of Figure 2 shows the evolution of the speed of light, normalized to its current value, over cosmic time. The central dashed line represents the best-fit values, the dot-dashed line denotes the null evolution, while the upper and lower solid lines represent the $1$-$\sigma$ confidence levels. The speed of light was up to $2.7$% slower at redshift $2$ compared to its current value. The right panel of Figure 2 shows the evolution of the Newtonian gravitational constant, normalized to its current value, over cosmic time. Like the speed of light, the central dashed line represents the best-fit values, and the upper and lower solid lines represent the 1-$\sigma$ confidence levels for the gravitational constant. The gravitational constant could have decreased to $10.4$% at $z=2$ compared to its current value, within the 1-$\sigma$ confidence levels. However, both constants include values consistent with the standard model values.
III.2.2 other bands with systematic error
Bands $b$-value $c(z=2)/c_{0}$ $G(z=2)/G_{0}$ SMC
$g$ $[-0.199\,,0.031]$ $[0.992\,,1.056]$ $[0.967\,,1.244]$ O
$r$ $[-0.067\,,0.035]$ $[0.990\,,1.019]$ $[0.962\,,1.076]$ O
$i$ $[-0.003\,,0.099]$ $[0.973\,,1.001]$ $[0.897\,,1.003]$ O
$z$ $[-0.039\,,0.063]$ $[0.983\,,1.011]$ $[0.933\,,1.044]$ O
all $[-0.057\,,0.033]$ $[0.991\,,1.016]$ $[0.964\,,1.065]$ O
Table 4: These are the upper bounds for the range of b obtained from each band, and the possible variations in the speed of light and Newton’s constant are derived from them when we include the systematic error. The current DES data shows results consistent with the standard model for all bands, including the $i$-band.
In Table 4, the results include systematic errors in the analysis of the parameter $b$, representing potential variations in the speed of light and Newton’s constant at redshift $z=2$. Here’s the detailed explanation for each band:
  • $g$ band: The $b$-value varies from $-0.199$ to $0.031$, meaning there is a broader range of possible deviations from the standard model. The speed of light ratio $c(z=2)/c_{0}$ lies between $0.992$ and $1.056$, allowing for slight variations where the speed could be either slower or faster than its current value. Newton’s constant $G(z=2)/G_{0}$ ranges from $0.967$ to $1.244$, suggesting possible deviations, but overall, the results are consistent with the SMC.
  • $r$ band: In this band, the $b$-value ranges from $-0.067$ to $0.035$. The speed of light ratio lies between $0.990$ and $1.019$, indicating only minor deviations in the speed of light at redshift $z=2$. The ratio of Newton’s constant ranges from $0.962$ to $1.076$, suggesting potential variations in Newton’s constant, but still consistent with the SMC.
  • $i$ band: The $b$-value ranges from $-0.003$ to $0.099$, providing upper and lower bounds on possible deviations. For the speed of light, its ratio can vary from $0.973$ to $1.001$, meaning the speed of light could have been slightly slower at redshift $z=2$. The ratio of Newton’s constant ranges from $0.897$ to $1.003$, implying that Newton’s constant could have been weaker. Despite the potential deviations, the i band supports the SMC when systematic errors are accounted for.
  • $z$ band: The $b$-value in the $z$ band ranges from $-0.039$ to $0.063$. The speed of light ratio lies between $0.983$ and $1.011$. The ratio $G(z=2)/G_{0}$ ranges from 0.933 to 1.044, suggesting only minor variations in Newton’s constant. The z band is consistent with the SMC.
  • All bands combined: When data from all bands are combined, the $b$-value ranges from $-0.057$ to $0.033$, providing a more constrained upper bound for potential variations. The speed of light ratio lies between $0.991$ and $1.016$, indicating only slight possible deviations. The ratio of Newton’s constant ranges from $0.964$ to $1.065$, showing limited potential variations in Newton’s constant. The combined data supports the SMC.
In summary, after including systematic errors, all bands, including the $i$ band, show consistency with the SMC. Table 4 demonstrates only slight potential variations in both the speed of light and Newton’s constant, suggesting that the current DES data generally supports the idea of constant fundamental constants at redshift $z=2$.
III.3 Systematic from z
@@
As can be seen from the above formula, when $b$ is positive (i.e., when the speed of light in the past is smaller than the current value and gradually increases), the value of $z_{\textrm{emit}}$ rises as the value of $b$ increases compared to the case where $b=0$ (i.e., when the speed of light is constant). Conversely, when $b$ is negative (i.e. when the speed of light in the past is larger than the current value and gradually decreases), the value of $z_{\textrm{emit}}$ decreases as the value of $b$ decreases.
$b$ $-0.05$ $-0.01$ 00 $0.01$ $0.05$
$z_{\textrm{emit}}$ 0.966 0.993 1 1.007 1.036
$\Delta z_{\textrm{emit}}$0.034 -0.007 0 0.007 0.036
$|\Delta z_{\textrm{emit}}|/z_{\textrm{emit}}(\%)$ 3.5 0.69 0 0.70 3.59
Table 5: These are changes in the redshift value obtained in the meVSL model depending on the value of $b$, compared to the known redshift value of $z=1$ in the SMC.
Table 5 presents how the redshift value ($z_{\textrm{emit}}$), its change ($\Delta z_{\textrm{emit}}$), and the percentage change ($|\Delta z_{\textrm{emit}}|/z_{\textrm{emit}}$) are influenced by the parameter $b$. As the value of $b$ varies from $-0.05$ to $0.05$ in increments of $0.01$, the redshift value also changes, ranging from $0.966$ to $1.036$. The difference in redshift is calculated by comparing each redshift value to the baseline redshift value at $b=0$. This change fluctuates between $-0.034$ and $0.036$. The last row shows the percentage change in redshift relative to the original $z$ value, with values ranging from $3.5$% to $3.59$%.
$b$ $-0.05$ $-0.01$ 00 $0.01$ $0.05$
$z_{\textrm{emit}}$ 1.921 1.984 2 2.017 2.086
$\Delta z_{\textrm{emit}}$0.079 -0.016 0 0.016 0.086
$|\Delta z_{\textrm{emit}}|/z_{\textrm{emit}}(\%)$ 3.97 0.82 0 0.83 4.31
Table 6: These are changes in the redshift value obtained in the meVSL model depending on the value of $b$, compared to the known redshift value of $z=2$ in the SMC.
We show how the redshift value ($z_{\textrm{emit}}$), its change ($\Delta z_{\textrm{emit}}$), and the percentage change ($|\Delta z_{\textrm{emit}}|/z_{\textrm{emit}}$) vary with the parameter $b$ in Table 6. The values of $b$ range from $-0.05$ to $0.05$. For each value of $b$, the corresponding redshift changes, ranging from $1.921$ to $2.086$. The difference in redshift is calculated by comparing each redshift value to the baseline value when $b=0$. The change in redshift fluctuates between $-0.079$ and $0.086$. We show the percentage of this change relative to the original redshift value, with the percentage ranging from $3.97$% to $4.31$% in the last row.
@@ IV Conclusion and Summary
Using the recently released dark energy survey supernova program’s $i$ band results, we applied the minimally extended varying speed of light model to investigate the cosmological evolution of the speed of light and Newton’s constant within the 1-$\sigma$ confidence levels. Our analysis indicates that the speed of light was between $0.4$% and $2.2$% slower at redshift $2$ compared to its current value. Furthermore, the gravitational constant may have decreased by $1.7$% to $8.4$% at $z=2$ compared to its present value. However, this result is obtained when we ignore the systematic effect due to the evolution of the stretch of supernova light curves in the $i$-band results. If we include this systematic error, we can obtain results for the speed of light or Newton’s constant that agree with the standard model even at the 1-sigma level.
These findings align with previous studies suggesting that Newton’s constant $G(z)$ decreasing with redshift could resolve the observed tension between the Planck15151515 best-fit $\Lambda$CDM cosmological background and the $\Lambda$CDM background favored by growth $f\sigma_{8}$ and weak lensing data. However, the full-band cosmological time dilation data from the dark energy survey supernova program matches the results from the standard model cosmology, indicating that these findings alone are not yet conclusive evidence of a varying speed of light. Additionally, high-z SNIa in the i-band may be affected by dust contamination or overwhelmed by noise, as indicated by slight deviations at wider widths.
It is important to note that the redshift $z$ values used in this data do not account for the additional effects predicted by the minimally extended varying speed of light model. For a consistent interpretation, the redshift values predicted by the minimally extended varying speed of light model should be used when analyzing the data. It necessitates recalculating the redshift for each dataset in the dark energy survey supernova program to compute the cosmological time dilation more accurately.
For this reason, we cannot yet definitively conclude that the Dark Energy Survey Supernova Program data rigorously demonstrates any changes in the speed of light. Nevertheless, through this manuscript, we aim to emphasize that the DES SN program data serves as a critical resource for testing variations in the speed of light and other physical constants over cosmic time. Furthermore, with future public releases of data that include a more rigorous analysis of redshift evolution, we hope to confirm whether any reliable evidence for time variation in physical constants exists.